OpenStudy (anonymous):

Find the x-intercept of the function f(x) = ax + 1 - 1. Answers available in ... 10 Take this time to do your best on this question. A) (0, 0) Eliminate Reactivate B) (a, 0) Eliminate Reactivate C) (-1, 0) Eliminate Reactivate D) (a - 1, 0) Eliminate Reactivate

Directrix (directrix):

Is f(x) = ax + 1 - 1 meant to be f(x) = a^x + 1 - 1? That would be f(x) = a^x when simplified.

OpenStudy (anonymous):

so wats da answer

Directrix (directrix):

I am trying to understand the question. Would you click on the blue DRAW and write the question? Thanks.

OpenStudy (anonymous):

wat blue draw

Directrix (directrix):

Look below for this:

OpenStudy (anonymous):